# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ 2 Confounding Ordinal Statistics and their Unobserved Gaps
If $\{X_{a},X_{b},\ldots,X_{[m]}\}$ is a random $m$-sample from a particular distribution, it will be important to know whether a factor $b$ follow or precede sequential changes in $a$ (e.g., events $a_{1},a_{2}$). For example, whether $a_{1}<b<a_{2}$ or $a_{1}<a_{2}<b$ lead to distinct scenarios in the estimation of $a$’s effect. The likelihood of such orders is determined by the DGP, as illustrated in Example.2. To that end, we define $X_{(1)},X_{(2)},\ldots,X_{(m)}$ to be a sorted version of the sample  [Gut, 2009], that is,
$X_{\min}\equiv X_{(1)}<X_{(2)}<\ldots\leq X_{(m)}\equiv X_{\max}$
@@
Notice that $X_{\min}$ is a random variable. For simplicity, we assume that, in sufficiently large samples and time resolution, the order values cannot coincide, leading to strict inequalities (as opposed to $\leq$).
@@
Suppose that $X_{1},X_{2},\ldots,X_{m}$ follow a continuous distribution with PDF $p$ and CDF $F$. We will use $X$, $U$ or $Z$ as alternative labels when considering distributions over observed, unobserved or both factors. Imagine a random sample of $U$ factors falling in such a way that its minimum is below a fixed observed value $x$. Fig.1(b) illustrates examples of draws where circles are samples of $U$ and interval ticks of $X$. The intervals show $X_{(1)}$ in samples with increasing dimensionality $m$. Circles to the left of the dashed line represent irreducible sample confounding effects. Under i.i.d. sampling, this can be expressed as
@@
which uses the observation that $x$ is the minimum if and only if all the $U$ are greater than $x$. The corresponding pdf is then $p_{X_{(1)}(x)}=m[1-F(x)]^{m-1}p(x)$.
In particular, we are interested in the distribution of the $k$-th sample gap and its associated variables,
@@
[…]
@@
where $\Delta_{(k)}m$ and $\Delta_{(k)}q$ are the number of observed and unobserved factors with positive probability in, respectively, an ordinal position or gap. These gap statistics are closely associated with the possible extensions among posets $\{G(X),G(Y)\}$ (such as those illustrated in Example.1).
An effect statistic for a gap will become relevant because it reflects only the effect of new causes introduced within the gap ($\Delta y(u)$, $a<u_{k}<b$ for two sample variables $a$ and $b$ and unobserved causes $u_{k}\subseteq U$). Of particular interest are gap effects when in-sample factors are fixed across ordinal observation sequences. Observed outcome differences in a gap $(k)$ reflect then exclusively the effect of unobserved factors falling in $(k)$. In a square diagram with $X=\{a,b,c\}$, for example, $Y_{(k)}(X{=}a)-Y_{(k+1)}(X{=}a)$ denotes outcome differences across its main diagonal, Fig.1(a). For the case of an increasing $m$, where $U_{(1)}\rightarrow\emptyset$, Eq.(5) suggest distinct levels of confounding across sample effect observations,
Proposition 1 (Partial-Observability and Effect Variance).
[…]
This suggests that even when a factor $b$ is not observed at a minimal order $(k=1)$, the unobserved confounding effects at $(k=1)$ can still be discounted by using other sample factor effects prior to $b$ in the order. For example, when $a$ is observed at $(k=1)$ but $b$ only at $(k=2)$, $\textrm{Var}[\Delta y_{(2)}(b\,|\,a)-\Delta y_{(1)}(a)]\leq\textrm{Var}[\Delta y_{(2)}(b\,|\,a)]$.
We will consider, in particular, the relationship between distributions of causes across gaps and (1) the number of permutations that $U{-}X$ systems can generate, and (2) their impact in the statistics of effect observations across sample ordinal positions. According to the separability principle (Principle.1), the first is a requirement for the separability of $U{-}X$ causes and confounders using effect observations. Some distributions of unobserved causes across gaps can asymptotically generate all distinct orders (all observable permutations of unobserved factors) and some can’t. This can be described by a multinomial coefficient, and by comparing the cycle structure of $U{-}X$ permutations to their ideal full set of permutations $q!$,
@@
We will study, in particular, $U{-}X$ systems where $\Delta_{(k)}m\gg\Delta_{(k)}q$ (fast) or $\Delta_{(k)}m\ll\Delta_{(k)}q$ (slow) in specific sampling periods. Fig.1(b, bottom-right) shows examples of this relationship in the former. The lower panel shows the limiting cases of $(\Delta m_{(1)},\Delta m_{(2)},...)=(m,m,...,)$ that can enumerate all orders (ratio $\nicefrac{{4!}}{{4!}}=1$), the case of (1,1,,)(1,1,...,)( 1 , 1 , … , ) that can enumerate only one ($\nicefrac{{1!}}{{4!}}=\nicefrac{{1}}{{4!}}$) and an intermediary case. The figure also shows the corresponding squares, and thus effect observations, available across the resulting samples.
@@
where $u_{0}\perp u_{\delta 0}$, $u_{1}\perp u_{\delta 1}$, etc. The first corresponds to the case of fixed (unobserved) effect backgrounds for a time of at least $m$, and it can lead to generalizeable effects in fast $U{-}X$ systems. The second correspond to cases where effects are observed at time delays $\delta$ (and assignment of unobserved confounders across gaps), $\pi_{\delta}$, no longer depends on the initial background, $\pi_{0}$ (Sect. 4 Related Work). It is useful in slow systems. Crucially, to formulate the $\delta$ required for background randomization, we will consider to what extent gap statistics allows us to identify linear extensions among posets $\{G(X),G(U)\}$, which, in turn, allow us to formulate $\delta$ theoretically, given the specific extension.
Confounding $U$ in samples following the conditions in Eq.(8) can be expressed in the form
@@
or, similarly, by saying that the variables $U_{(k)}$ form an (additive) Markov chain. The observation that order statistics under the previous conditions form a Markov chain was first made by Kolmogorov  [Rényi, 1953]. If $U$ follow both exponential and uniform distributions gap effects will have this additive property. While this not always the case in generic $U{-}X$ non-experimental samples, decomposition of samples in subsets of effect observations with these properties will offer comparable advantages. Eq.(9) suggests that confoudness of effects in squares increases linearly across their columns, and is trivially associated with an ANOVA variance decomposition for effect variance in samples (Appendix G).
Figure 2: (a) sampling sequences across sample, $X=\{a,b,c\}$, gaps and confounder partition $U=\{\{u_{1}\},\{u_{2},u3\},\{u_{4}\}\}$, (d) normalized variation separation distance (distance to a Uniform distribution) of all backgrounds $\pi_{0}$ to other backgrounds after time $t$, vertical dashed lines mark limits on successive background randomness based on their inversions (top panel shows the same for hamming distances across all sample effect observations), (c) randomization in graphical models of in-sample treatment $A$, other factors $W$, unobserved factors $U$, and outcome $Y$, (d) combinatorial relations in a square ($m=2$), circles are differences and gray circles intersections of indicated letters between $x_{0}$ and other units.
@@ 3 Sampling Treatment-Background Assignments
We will start with a definition for generalizeable sample effect sequences, as permutations of remainders of $Z/X=U$ for a fixed $X$ with unobserved factors $U$. This will formulate what sampling sequences, and thus subsets of permutations $\Pi_{n}(Z/X)$, are necessary to make claims about the generalizability of effects in $X$. Eq.(3) formulated effects for a single fixed factor $a$, where backgrounds were $\Pi_{n}(Z/\{a\})$, and is generalized here for samples $X$. The next results rely on the following assumptions, which are somewhat typical,
@@
    Constant DGP. The set of observed, $X$, and unobserved factors, $U$, in $Z=X\cup U$ do not change during the sampling period. This implies no measurement errors, and that, conditioned on all other factors in $Z$, probabilities and effect of causes do not change or have low variance. This leads to typical notions of irreducible variance and errors in samples  [Tibshirani et al., 2001].
@@
    Multiple Exposures. Effects remain constant under multiple exposures. While this is implied by most definitions of effects, it is inadequate for continuous and cumulative effect interventions, such as treatments involving dosages and other interactions with the numbers of exposures.
@@
    Large-scale data. The sample $X$ includes both causes and spurious factors. In our cases of interest, such large-scale economic, genetic and medical data these are typically available.
@@
    Non-strongly antagonistic interactions. Principle.(1) imply additive effects among causes (see Appendix.B for a proof). This is a weaker version of an assumption of additivity of unobserved causes’ effects, useful in samples that do contain causes in such ideal conditions [^1] Many times, true causes are unobserved but we still want to find the closest surrogate in-sample factor that reproduces the cause effects across observable conditions. In these cases, additivity or monotonicity of effects might be too strong a requirement, and other non-interaction requirements for unobserved causes and gaps can be formulated. These requirements are unnecessary for Principle.1, except for strongly antagonistic effects, and unnecessary in the empirical cases we have considered so far. Interaction effects for in-sample effects are adjusted by averaging across square in-sample backgrounds, Eq.(3). .
Imagine our effect observations to be discrete and equally spaced points in time. If an effect sample at time $t$ is next sampled at time $t^{\prime}$, we write $b(t)=t^{\prime}$. If $t$ is an integer then let $t\bmod q$ denote the remainder when $t$ is divided by $q$. Finally we let $\bar{q}$ be a shorthand for the set of possible remainders $\{0,1,2,...,q-1\}$. Sampling sequences can be interpreted mathematically in several ways. We choose to think of them as permutations $\pi:b(t)$ of the set $K$ of integer sample gap indices. We call each factor $k$ in $\bar{q}$ an individual sample gap index, $u_{k}$, in reference to Eq.(2), and we call a permutation of $\bar{q}$ a single background. In respect to sample effects, each permutation defines, in turn, a distinct treatment-background assignment. Regularities in this assignment in $U{-}X$ systems defines their treatment-background mechanisms. Thus a sampling strategy defines a function $b:K\rightarrow K$, where $b(t)$ is the next sample gap to be observed at time $t\in\mathbb{N^{+}}$. Remember permutations can have fixation points. Each such effect observation correspond to a subset of unobserved factors $u_{k}\subseteq U$, Eq.(2), and we say the effect of factor $a$ is observed in the background of $u_{k}$ in this case.
@@
Fig.2(a) shows a sample $X_{3}=\{a,b,c\}$ with a fixed order statistic, $a<b<c$, and changing $U_{4}=\{u_{1},u_{2},u_{3},u_{4}\}$ (slow sampling). The diagram shows the transition of backgrounds under repeated observations of effects with a fixed-order $a<b<c$. Backgrounds are partitioned in this example as $\{\{u_{1}\},\{u_{2},u_{3}\},u_{4}\}$. Each arrow correspond to the sequential mapping of these unobserved factors (arrow labels) across sample gaps (intervals).
Effect backgrounds $Z/X$ are not necessarily constant due to the distinct frequencies, correlations and periodicities among $U$ and $X$ factors. A key definition in this article is
@@
A finite sampling sequence $b$ of nonnegative integers $b(0)...b(q-1)$, and its $U{-}X$ assignments, is EV-increasing if and only if the integers $[t+b(t)]_{q}$ are distinct.
In Fig.2(a), for example, the top and bottom diagrams illustrate sequences that are not EV increasing, as each variable effect is observed in only one background. The middle diagram is EV increasing with period of $2$, as effects are observed in new backgrounds at each time, and all enumerable backgrounds for a fixed $Z/X$ are observed after time $2$. A complementary property for sampling sequences is
@@
In Fig.2(a), the top and middle diagrams are CF non-increasing. The bottom is CF non-increasing for a period of $2$. The ***-marked extensions in Example.1 are CF non-increasing, but so are any sequences with cyclic permutations of in-sample factors under the condition that the unobserved background is fixed (i.e., subsequent rows of sample squares).
@@
This leads to the notion of enumeration of effect backgrounds,

### v5

@@ 2 Ordinal Statistics and Sample Confounding Gaps
If $\{X_{a},X_{b},\ldots,X_{m}\}$ are random samples from a particular DGP distribution on $X_{m}$, it will be important to know whether changes in a factor $a$ are expected to follow or precede changes in another factor $b$. The likelihood of such order relations is determined by the DGP, as illustrated in Example. (2). To describe such patterns more formally, we define $X_{(1)},X_{(2)},\ldots,X_{(m)}$ to be an ordered version of sample factors  [Gut, 2009], that is,
$X_{\min}\equiv X_{(1)}<X_{(2)}<\ldots<X_{(m)}\equiv X_{\max}$
@@
Notice that $X_{\min}$ is a random variable. For simplicity, we assume that, in sufficiently large samples and time resolution, ranks do not coincide, leading to strict inequalities (as opposed to $\leq$).
@@
Suppose that $X_{1},X_{2},\ldots,X_{m}$ follow a continuous distribution with PDF $p$ and CDF $F$. We will use $X$, $U$ or $Z$ as alternative labels when considering distributions over observed, unobserved or either factors. Imagine, in particular, a random sample of $U$ factors when their minimum is below an in-sample value $x$. Fig.1(b) give examples where circles are $U$ samples and interval ticks are $X$ samples. The successive lines show $X_{(k)}$ in samples with increasing dimensionality $m$. Circles to the left of the dashed line, and $X_{(1)}$, represent irreducible sample confounding effects. Under i.i.d. sampling, this can be expressed as
@@
which uses the observation that $x$ is the minimum if and only if all $X$ are greater than $x$ [Gut, 2009].
We are specially interested in the distribution of the $k$-th sample gap and its associated variables,
@@
where $u\in\mathcal{P}(U)$. In this notation, $\Delta X_{(k)}$ and $\Delta y_{(k)}$ are gap and gap-effect distributions, and $\Delta U_{(k)}$ is the distribution of unobserved confounders across gaps. We also define the counts
@@
where $\Delta_{(k)}m$ and $\Delta_{(k)}q$ are the number of, respectively, observed and unobserved factors with positive probability in the sample’s $k$-th ordinal position or gap. These gap statistics are closely associated with the possible extensions among posets $\{G(X),G(Y)\}$ (such as those illustrated in Example. (1)).
Sample gap effect statistics will become relevant, in particular, because they reflect only the effect of new causes introduced within the gap ($\Delta y(u)$, $a<u_{k}<b$ for two sample factors $a$ and $b$ and unobserved causes $u_{k}\subseteq U$). Of particular interest are gap effects when in-sample factors are the same in subsequent effect observations. In this case, outcome differences in a gap $k$ reflect exclusively the effect of unobserved factors ’falling’ in $k$, $0<k\leq m$. In a square diagram with $X=\{a,b,c\}$, for example, $Y_{(k)}(X{=}a)-Y_{(k+1)}(X{=}a)$ denotes differences in outcome across its main diagonal, Fig.1(a). Eq.(5) suggest distinct levels of confounding across a sample’s effect observations,
Proposition 1 (Sample Ordinal Statistics and Effect Variance).
Effect observations of factor $a$ at $\textrm{argmin}_{k}X_{(k)}=\{a\}$ are a sample’s minimally confounded and maximally precise individual observations of $a$’s effect, $\textrm{Var}^{-1}[\Delta y(a)]$. In a single-row square, the effect of its first factor is observed at maximum precision. All other effects can be estimated confounded only by $\Delta_{(k)}y$, by conditioning on effect observations prior in the square row order (and their interaction effects).
Proposition 2 (Full-Observability and Effect Variance).
In a square with $m$ factors and $q=0$, effects of all factors $X_{m}$ can be observed, simultaneously, at maximum precision and unconditionally.
This suggests that, even when a factor $b$ is not observed at a minimal order $k=1$, unobserved confounding effects at $k=1$ can still be discounted from estimates by using other sample factor effects prior to $b$ in row order. For example, when $a$ is observed at $k=1$ but $b$ only at $k=2$, $\textrm{Var}[\Delta y_{(2)}(b\,|\,a)-\Delta y_{(1)}(a)]\leq\textrm{Var}[\Delta y_{(2)}(b\,|\,a)]$.
We will consider, in particular, the relationship between the number of permutations that $U{-}X$ systems can generate and their impact in the statistics of effects across sample gaps. According to Principle (1) the first is a requirement for the separability of causes and confounders, when using effect observations. Some distributions of unobserved causes across gaps can asymptotically generate all distinct orders (all permutations of unobserved factors) and some can’t. This can be described by a multinomial coefficient, or, by dividing the set of possible permutations in a $U{-}X$ system, $Pi_{n}(Z/\{a\})$, and the ideal full set of permutations $q!$,
@@
This can be defined similarly for in-sample factors $X$. Fig.1(b, bottom and right) shows examples. The bottom panel shows the limiting cases of $(\Delta m_{(1)},\Delta m_{(2)},...)=(m,m,...,)$ that can enumerate all orders (ratio $\nicefrac{{4!}}{{4!}}=1$), the case of (1,1,,)(1,1,...,)( 1 , 1 , … , ) that can enumerate only one ($\nicefrac{{1!}}{{4!}}=\nicefrac{{1}}{{24}}$) and an intermediary case. The figure also shows the corresponding squares (right), and thus effect observations, available across the resulting samples.
@@
where $u_{0}\perp u_{\delta 0}$, $u_{1}\perp u_{\delta 1}$, etc. The first corresponds to the case of fixed (unobserved) effect backgrounds for a time period of at least $m$, and it can lead to generalizeable effects in fast $U{-}X$ systems. The second correspond to cases where effects are observed at time delays $\delta$, when assignment of unobserved confounders across gaps, $\pi_{\delta}$, no longer depends on the initial background, $\pi_{0}$. It is useful in in-sync and slow systems.
Confounding in samples following the conditions in Eq.(8) can be expressed in the form
@@
or, similarly, by saying that the variables $\Delta U_{(k)}$ form an (additive) Markov chain. The observation that order statistics under the previous conditions form a Markov chain was first made by Kolmogorov  [Rényi, 1953]. If $U$ follows both exponential and uniform distributions, gap effects will have this additive property. While this is not always the case in non-experimental samples, decomposition of samples in subsets of effect observations with these properties will offer some advantages, in particular, in respect to sample error decomposition. More specifically than Proposition (1), Eq.(9) suggests that confoudness of effects in a square increases linearly across its columns, and is trivially associated with an ANOVA variance decomposition for effect variance in samples (Appendix G).
Figure 2: (a) sampling sequences in a sample, $X=\{a,b,c\}$, with confounder partition $U=\{\{u_{1}\},\{u_{2},u3\},\{u_{4}\}\}$, (b) normalized variation separation distance (distance to a Uniform distribution) of all backgrounds $\pi_{0}$ to other backgrounds after time $t$, vertical dashed lines mark limits on successive background randomness based on their inversions, Eq.(11) (top panel shows the same without a test of stationarity on gap distributions $p$), (c) randomization in graphical models with in-sample treatment $A$, other factors $W$, unobserved factors $U$, and outcome $Y$, (d) combinatorial relations in a square ($m=2$), circles are differences and gray circles are inter-unit factor value intersections (letters) between $x_{0}$ and other units.
@@ 3 Sampling Treatment-Background Assignments
We will start with a definition for sampling sequences, and subsets of permutations $\Pi_{n}(Z/X)$, when we can make repeated measurements for the observed variables, $X$, in the DGP. Eq.(3) formulated a definition of effects for a single fixed factor $a$, where backgrounds were $\Pi_{n}(Z/\{a\})$. It is generalized here for samples $X$. These definitions rely on the following assumptions, which are somewhat typical,
@@
    Constant DGP. The set of observed, $X$, and unobserved factors, $U$, in $Z=X\cup U$ do not change during the sampling period and there are no measurement errors. This leads to typical notions of irreducible variance and errors in samples  [Tibshirani et al., 2001].
@@
    Multiple Exposures. Effects remain constant under multiple exposures. While often assumed, this might be inadequate for interventions with continuous and cumulative effects, such as treatments involving dosages or complex interactions with exposures.
@@
    Large-scale data. The sample $X$ includes both causes and spurious factors. In our applications of interest, such as large-scale economic, genetic and medical data, samples typically contain both.
@@
    Non-strongly antagonistic unobserved interactions. When samples contain true causes Principle (1) holds (Appendix.B). Many times, however, causes are unobserved but we still want to find in-sample surrogates that closely reproduce their root cause effects. In this case, Principle (1) might not hold when the effects of unobserved causes (i.e., observed in the same sample gap) strongly interact and can cancel each other.
Sampling sequences can be interpreted mathematically in several ways. We choose to think of them as permutations $\pi:b(t)$ of the set $K$ of integer sample gap indices. Imagine our effect observations to be discrete and equally spaced points in time. If an effect, sampled at time $t$, is next sampled at time $t^{\prime}$, we write $b(t)=t^{\prime}$. If $t$ is an integer then let $t\bmod q$ denote the remainder when $t$ is divided by $q$. Finally, we let $[t]_{q}$ be a shorthand for the set of possible remainders $\{0,1,2,...,q-1\}$ in this division. We call each factor $k$ in $[t]_{q}$ an individual sample gap index, $u_{k}$, Eq.(2). Remember that permutations can have fixation points (positions whose order do not change). Each individual sample effect observation is thus associated with a subset of unobserved factors $u_{k}\subseteq U$, Eq.(2), and we say, in this case, that the effect is observed in the background of $u_{k}$. Considering multiple sample factors, each permutation defines, in turn, distinct treatment-background assignments. Regularities across assignments in $U{-}X$ systems define their treatment-background mechanism.
@@
Fig.2(a) illustrate this concept in a sample $X_{3}=\{a,b,c\}$ with a fixed order statistic, $a<b<c$, and changing $U_{4}=\{u_{1},u_{2},u_{3},u_{4}\}$ (slow sampling). The diagram shows the transition of backgrounds under repeated observations of effects with a fixed-order $a<b<c$. Backgrounds are partitioned in this example as $\{\{u_{1}\},\{u_{2},u_{3}\},u_{4}\}$. Each arrow indicates the sequential mapping of unobserved factors (arrow labels) across sample gaps (intervals). See  [Buhler and Graham., c2004.] for analogous combinatorial concepts.
Effect backgrounds $Z/X$ are not necessarily constant due to distinct frequencies, correlations and periodicities among $U$ and $X$ factors. A key definition in this article is
@@
A finite sampling sequence $b$ of nonnegative integers $b(0)...b(q-1)$, and its $U{-}X$ assignments, are EV-increasing if and only if the integers $[t+b(t)]_{q}$ are distinct.
In Fig.2(a), for example, the top and bottom diagrams illustrate sequences that are not EV increasing, as effects for each variable effect are observed in only one background. The middle diagram is EV increasing with period $2$, as effects are observed in new backgrounds for that period, and all enumerable backgrounds for a fixed $Z/X$ are observed after time $2$. A complementary property for sampling sequences is
@@
In Fig.2(a), the top and middle diagrams are CF non-increasing. The bottom is CF non-increasing for a period of $2$. In Example (1), the sequences $\mathbf{ab}u_{1}u_{2}u_{3}\,,\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,,u_{1}u_{2}\mathbf{ab}u_{3},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab}.$ are CF non-increasing, but so are any sequences with cyclic permutations of in-sample factors (i.e., subsequent rows of sample squares) under the condition that the unobserved background is fixed.
@@
These sampling strategies can be described with modular arithmetic, $b(t+1)=c_{1}\times b(t)+c_{2}\bmod q$, with, respectively, multiplicative ($c_{1}\gg 1$) or additive ($c_{2}\gg 0$) constants. In abstract algebra, a group is a mathematical structure consisting of a set of elements and an operation (like addition or multiplication) that satisfies certain properties. In modular arithmetic, the integers coprime (relatively prime) to $q$ from the set $\{0,1,\dots,q-1\}$ of $q$ non-negative integers form a group under multiplication, called the multiplicative group of integers modulo $q$. Equivalently, the elements of this group can be thought of as the congruence classes, ’residues’ modulo $q$, that are coprime to $q$.
Sequences generated by recursive multiplication ($c_{1}\gg 1$) exhibits more non-linear and less predictable behavior compared to addition ($c_{1}\approx 1$). One group-theoretic property of interest is closure, which states that the result of performing the operation on any two elements of the group is still an element of the group. In the context of addition modulo $q$, addition satisfies the closure property. This means that for any two residues $i$ and $j$ modulo $q$, $i+j$ is still a residue modulo $q$. However, when we consider multiplication modulo $q$, closure is not guaranteed. Not every pair of residues modulo $q$ has a product that remains within the set of residues modulo $q$. This lack of closure allows multiplicative recursions ($c_{1}\gg 1$) to include numbers in different residue classes modulo $q$, and, at the same time, introduce more complexity and non-linearity compared to addition. This observation is often employed by random number generators  [Knuth, 1997] as means to generate random numbers.
3.1 Background Enumeration
When $q\gg m$, the constant $c_{1}$ describes the rate of increase in the counts of out-sample factors, and $c_{2}$ of in-sample factors. We first consider fast sampling, the case where $\Delta_{(k)}m>\Delta_{(k)}q$ in specific sampling periods. In this case, in-sample effects are observed for all factors in constant backgrounds. The constant $c_{2}$ describe ’phase’ differences that, when $c_{2}$ and $q$ are co-prime, generate samples with the cyclic permutations of $X$ (individual square rows). Systems under fast sampling are EV-increasing for a constant sampling rate when $c_{1}/q$ is integral,